(sinx+cosx)2 -sin(2x)=1


1
Expert's answer
2021-05-07T14:31:18-0400

"(\sin{x}+\cos{x})^2-\sin{2x}=1\\
\sin^2{x}+\cos^2{x}+2\sin{x}*\cos{x}-\sin{2x}=1\\
\sin^2{x}+\cos^2{x}+2\sin{x}*\cos{x}-(2\sin{x}*\cos{x})=1\\
1+2\sin{x}*\cos{x}-(2\sin{x}*\cos{x})=1\\
1=1"

Answer:"x \in (-\infty;\infty)"